Aiden paid $30 to be a member of the Fox Lake Gym. When he takes the boxing class, it costs him$2. Noah is not a member of the Gym, and it costs him $5 for the same class. How many classes would it take for Aiden's total cost to equal Noah's total cost?

Found 2 solutions by math_helper, MathTherapy:
Answer by math_helper(2461)


What you need to do is figure out Aiden's (A) and Noah's (N) costs a function of number of classes taken:
Let x = number of boxing classes
A = 30 + 2x
N = 5x
Their costs are equal when A = N, so set the two equations equal to each other and then solve for x:
30 + 3x = 5x
30 = 2x
x = 30/2 = 15
When they've taken 15 boxing classes, their costs are the same. Beyond 15 classes, Aiden's cost will be LESS than Noah's, and for any number less than 15 classes, Noah's cost will be LESS than Aiden's.

Answer by MathTherapy(10806)


Aiden paid $30 to be a member of the Fox Lake Gym. When he takes the boxing class, it costs him$2. Noah is not a member of the Gym, and it costs him $5 for the same class. How many classes would it take for Aiden's total cost to equal Noah's total cost?
Mathhelper made a mistake. 
It actually takes 10 classes for their total cost to be the same.